OpenStudy (anonymous):

write the equation of the line parellel to y=5x-7 through the point (4,1)

OpenStudy (anonymous):

the slope has to be 5 (to be parallel) y = 5x + b to solve for b plug in the point: 1 = 5(4) + b 1 = 20 + b b = -19 so the equation should be: y = 5x - 19